# thanks: Corresponding Author

### v1

@@ III.1 $\mathcal{M}\rightarrow\eta_{b}\eta_{c}$
×$\times J^{\eta_{c}}(0)J^{\dagger}(x)\}|0\rangle,$
@@
×$\times\langle\eta_{b}(p^{\prime})\eta_{c}(q)|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
where $m_{\eta_{b}}=(9398.7\pm 2.0)~{}\mathrm{MeV}$ and $m_{\eta_{c}}=(2983.9\pm 0.4)~{}\mathrm{MeV}$ are the masses of the quarkonia $\eta_{b}$ and $\eta_{c}$ PDG:2024 , respectively. The correlation function $\Pi_{1}^{\mathrm{Phys}}(p,p^{\prime})$ is a contribution of the ground-state particles, whereas the dots denote effects of higher resonances and continuum states. Next, we introduce the matrix elements
@@
where $f_{\eta_{b}}=724~{}\mathrm{MeV}$ and $f_{\eta_{c}}=(421\pm 35)~{}\mathrm{MeV}$ are decay constants of the mesons $\eta_{b}$ and $\eta_{c}$, respectively.
@@
×$\times\frac{m^{2}+m_{\eta_{b}}^{2}-q^{2}}{\left(p^{\prime 2}-m_{\eta_{b}}^{2}\right)(q^{2}-m_{\eta_{c}}^{2})}+\cdots.$ (20)
@@
The same correlation function $\Pi_{1}(p,p^{\prime})$ computed in terms of the heavy quark propagators is
@@
×$\left.\times\gamma_{5}S_{c}^{bj}(x)\gamma_{5}S_{c}^{ja}(-x)\gamma_{5}S_{b}^{ai}(x-y)\right].$ (21)
Due to a triviality of its structure $\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})$ is equal to the amplitude $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$. This amplitude after the double Borel transformations over $-p^{2}$, $-p^{\prime 2}$ and continuum subtractions determines the SR for $g_{1}(q^{2})$
@@
×$\times e^{m^{2}/M_{1}^{2}}e^{m_{\eta_{b}}^{2}/M_{2}^{2}}\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (22)
where
[…]
×$\times e^{-s/M_{1}^{2}}e^{-s^{\prime}/M_{2}^{2}}.$ (23)
It is written using the spectral density $\rho_{1}(s,s^{\prime},q^{2})$ containing the perturbative and dimension-$4$ components.
The sum rule in Eq. (22) contains as auxiliary quantities the Borel $\mathbf{M}^{2}=(M_{1}^{2},M_{2}^{2})$ and continuum threshold parameters $\mathbf{s}_{0}=(s_{0},s_{0}^{\prime})$. To perform numerical analysis, we employ in the $\mathcal{M}$ channel $M_{1}^{2}$ and $s_{0}$ from Eq. (LABEL:eq:Wind1). In the $\eta_{b}$ channel as the parameters $(M_{2}^{2},\ s_{0}^{\prime})$ we use
@@
A part ±$\pm 13.7$ of the total errors in Eq. (29) is connected with the uncertainties of the coupling $g_{1}$, whereas ±$\pm 6.1$ generated by the ambiguities of the particles’ masses in Eq. (27).
@@ III.2 $\mathcal{M}\rightarrow J/\psi\Upsilon$
$\Pi_{\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{\Upsilon}(y)$ (30)
×$\times J_{\nu}^{J/\psi}(0)J^{\dagger}(x)\}|0\rangle,$
@@
To find the phenomenological component $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ of the sum rule, we need to rewrite Eq. (30) using the involved particles’ physical parameters. By taking into account only contributions of the ground-level states, we recast the correlator $\Pi_{\mu\nu}(p,p^{\prime})$ into the form
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{\Upsilon}|\Upsilon(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{\Upsilon}^{2}}\frac{\langle 0|J_{\nu}^{J/\psi}|J/\psi(q,\varepsilon_{2})\rangle}{q^{2}-m_{J/\psi}^{2}}$
×$\times\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}(p\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
Besides, one should specify the matrix element $\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}(p)\rangle$ which depends on the momenta and polarization tensor and vectors of the particles $\mathcal{M}$, $\Upsilon$ and $J/\psi$
@@
×$\left.\times\varepsilon_{1}^{\ast}\cdot\varepsilon_{2}^{\ast}-q\cdot\varepsilon_{1}^{\ast}p^{\prime}\cdot\varepsilon_{2}^{\ast}\right].$ (34)
As a result, for $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ we get the expression
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=g_{2}(q^{2})\frac{\Lambda f_{\Upsilon}m_{\Upsilon}f_{J/\psi}m_{J/\psi}}{\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{\Upsilon}^{2})(q^{2}-m_{J/\psi}^{2})}$
×$\times\left[\frac{(m^{2}-m_{\Upsilon}^{2}-q^{2})}{2}g_{\mu\nu}-p_{\nu}^{\prime}q_{\mu}\right]+\cdots.$ (35)
The function $\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$, i.e., the QCD component of the sum rule is given by the expression
$\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ib}(y-x)\right.$
×$\left.\times\gamma_{5}S_{c}^{bj}(x)\gamma_{\nu}S_{c}^{ja}(-x)\gamma_{5}S_{b}^{ai}(x-y)\right].$ (36)
@@
×$\times e^{m^{2}/M_{1}^{2}}e^{m_{\Upsilon}^{2}/M_{2}^{2}}\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$
@@
for $(M_{2}^{2},s_{0}^{\prime})$ satisfy all these requirements. Note that $s_{0}^{\prime}$ is limited by the mass $m_{\Upsilon(2S)}=(10023.4\pm 0.5)~{}\mathrm{MeV}$of the radially excited state $\Upsilon(2S)$, i.e., $s_{0}^{\prime}<m_{\Upsilon(2S)}^{2}$.
@@ III.3 $\mathcal{M}\rightarrow B_{c}^{+}B_{c}^{-}$
×$\times J^{B_{c}^{-}}(0)J^{\dagger}(x)\}|0\rangle,$ (42)
@@
$\langle B_{c}^{+}(p^{\prime})B_{c}^{-}(q)|\mathcal{M}\rangle=g_{2}(q^{2})p\cdot p^{\prime},$ (45)
@@
×$\times\frac{m^{2}+m_{B_{c}}^{2}-q^{2}}{2(q^{2}-m_{B_{c}}^{2})}+\cdots.$ (46)
@@
×$\left.\times\gamma_{5}S_{b}^{ai}(x-y)\right]\mathrm{Tr}\left[\gamma_{5}S_{c}^{bj}(x)\gamma_{5}S_{b}^{jb}(-x)\right].$ (47)
@@
×$\times e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}}^{2}/M_{2}^{2}}\Pi_{3}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (48)
@@ IV.1 $\mathcal{M}\rightarrow D^{0}\overline{D}^{0}$ and $D^{+}D^{-}$
$\widetilde{\Pi}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{D^{0}}(y)$ (52)
×$\times J^{\overline{D}^{0}}(0)J^{\dagger}(x)\}|0\rangle.$
@@
First, we find the physical side of the SR and write down $\widetilde{\Pi}(p,p^{\prime})$ in the following way
$\widetilde{\Pi}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J^{D^{0}}|D^{0}(p^{\prime})\rangle}{p^{\prime 2}-m_{D^{0}}^{2}}\frac{\langle 0|J^{\overline{D}^{0}}|\overline{D}^{0}(q)\rangle}{q^{2}-m_{D^{0}}^{2}}$
×$\times\langle D^{0}(p^{\prime})\overline{D}^{0}(q)|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
To simplify $\widetilde{\Pi}^{\mathrm{Phys}}(p,p^{\prime})$, we rewrite it in terms of the involved particles’ masses and current couplings (decay constants). The matrix element of the pseudoscalar $D^{0}$ ($\overline{D}^{0}$) meson is given by the expression
@@
Afterwards, we recast $\widetilde{\Pi}^{\mathrm{Phys}}(p,p^{\prime})$ into the following form
$\widetilde{\Pi}^{\mathrm{Phys}}(p,p^{\prime})=\widetilde{g}_{1}(q^{2})\frac{\Lambda f_{D}^{2}m_{D}^{4}}{2m_{c}^{2}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{D^{0}}^{2}\right)}$
×$\times\frac{\left(m^{2}+m_{D^{0}}^{2}-q^{2}\right)}{(q^{2}-m_{D^{0}}^{2})}+\cdots,$ (57)
Note that the whole expression in Eq. (57) is the invariant amplitude $\widetilde{\Pi}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ which can be employed to get SR for the form factor $\widetilde{g}_{1}(q^{2})$.
@@
$\widetilde{\Pi}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{b}b\rangle$
×$\times\mathrm{Tr}\left[\gamma_{5}S_{c}^{ja}(y-x){}S_{c}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right],$ (58)
@@
The correlation function $\widetilde{\Pi}^{\mathrm{OPE}}(p,p^{\prime})$ depends on three propagators and vacuum condensate $\langle\overline{b}b\rangle$ of $b$ quarks. This function differs from a standard correlators which contain four heavy quark propagators $S_{b(c)}(x)$. In fact, to calculate $\Pi^{\mathrm{OPE}}(p,p^{\prime})$ one contracts heavy and light quark fields, and because  mesons $D^{0}\overline{D}^{0}$ do not contain $b$ quarks gets an expression with the condensate $\langle\overline{b}b\rangle$.
Having activated an equality that relates heavy quark and gluon condensates
$\langle\overline{b}b\rangle=-\frac{1}{12m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle,$ (59)
we get
$\widetilde{\Pi}^{\mathrm{OPE}}(p,p^{\prime})=-\frac{1}{36m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
×$\times\mathrm{Tr}\left[\gamma_{5}S_{c}^{ja}(y-x){}S_{c}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (60)
Stated differently, the correlator $\widetilde{\Pi}^{\mathrm{OPE}}(p,p^{\prime})$ is suppressed by the factor $\langle\alpha_{s}G^{2}/\pi\rangle$. We denote the corresponding invariant amplitude as $\widetilde{\Pi}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ and carry out all required operations to fix SR for the form factor $\widetilde{g}_{1}(q^{2})$.
We omit standard details and present only final results for the process $\mathcal{M}\rightarrow D^{0}\overline{D}^{0}$. Numerical computations of the form factor $\widetilde{g}_{1}(q^{2})$ is performed using

### v2

@@ III.1 $\mathcal{M}\rightarrow\eta_{b}\eta_{c}$
$\times J^{\eta_{c}}(0)J^{\dagger}(x)\}|0\rangle,$
@@
$\times\langle\eta_{b}(p^{\prime})\eta_{c}(q)|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
where $m_{\eta_{b}}=(9398.7\pm 2.0)~\mathrm{MeV}$ and $m_{\eta_{c}}=(2983.9\pm 0.4)~\mathrm{MeV}$ are the masses of the quarkonia $\eta_{b}$ and $\eta_{c}$ PDG:2024 , respectively. The term in the correlation function $\Pi_{1}^{\mathrm{Phys}}(p,p^{\prime})$ is a contribution of the ground-state particles, whereas the dots denote effects of higher resonances and continuum states. Next, we introduce the matrix elements
@@
where $f_{\eta_{b}}=724~\mathrm{MeV}$ and $f_{\eta_{c}}=(421\pm 35)~\mathrm{MeV}$ are decay constants of the mesons $\eta_{b}$ and $\eta_{c}$.
@@
$\times\frac{m^{2}+m_{\eta_{b}}^{2}-q^{2}}{\left(p^{\prime 2}-m_{\eta_{b}}^{2}\right)(q^{2}-m_{\eta_{c}}^{2})}+\cdots.$ (20)
@@
The same correlation function $\Pi_{1}(p,p^{\prime})$ computed in terms of the heavy quark propagators reads
@@
$\left.\times\gamma_{5}S_{c}^{bj}(x)\gamma_{5}S_{c}^{ja}(-x)\gamma_{5}S_{b}^{ai}(x-y)\right].$ (21)
Due to a triviality of its structure $\Pi_{1}^{\mathrm{OPE}}(p,p^{\prime})$ is equal to the amplitude $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$. Having equated the amplitudes $\Pi_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$, performed the double Borel transformations over $p^{2}$, $p^{\prime 2}$ and continuum subtractions, one gets the sum rule for $g_{1}(q^{2})$
@@
$\times e^{m^{2}/M_{1}^{2}}e^{m_{\eta_{b}}^{2}/M_{2}^{2}}\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (22)
In Eq. (22) the correlator $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ has the following form
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{4(m_{b}+m_{c})^{2}}^{s_{0}}ds\int_{4m_{b}^{2}}^{s_{0}^{\prime}}ds^{\prime}\rho_{1}(s,s^{\prime},q^{2})$
$\times e^{-s/M_{1}^{2}}e^{-s^{\prime}/M_{2}^{2}}.$ (23)
It is written using the spectral density $\rho_{1}(s,s^{\prime},q^{2})$ which contains the perturbative and dimension-$4$ components. The function $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ depends on the Borel $\mathbf{M}^{2}=(M_{1}^{2},M_{2}^{2})$ and continuum threshold parameters $\mathbf{s}_{0}=(s_{0},s_{0}^{\prime})$. The pair of parameters $(M_{1}^{2},s_{0})$ corresponds to the hadronic $\mathcal{M}$ molecule channel, whereas $(M_{2}^{2},s_{0}^{\prime})$ describes the channel of the $\eta_{b}$ meson.
To perform numerical analysis, in the $\mathcal{M}$ channel we employ $M_{1}^{2}$ and $s_{0}$ from Eq. (LABEL:eq:Wind1). In the $\eta_{b}$ channel as the parameters $(M_{2}^{2},\ s_{0}^{\prime})$ we use
@@
A part $\pm 13.7$ of the total errors in Eq. (29) is connected with the uncertainties of the coupling $g_{1}$, whereas $\pm 6.1$ generated by the ambiguities of the particles’ masses in Eq. (27).
@@ III.2 $\mathcal{M}\rightarrow J/\psi\Upsilon$
$\Pi_{2\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{\Upsilon}(y)$ (30)
$\times J_{\nu}^{J/\psi}(0)J^{\dagger}(x)\}|0\rangle,$
@@
To find the phenomenological component $\Pi_{2\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ of the sum rule, we need to rewrite Eq. (30) using the involved particles’ physical parameters. By taking into account only contributions of the ground-level states, we recast the correlator $\Pi_{2\mu\nu}(p,p^{\prime})$ into the form
$\Pi_{2\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{\Upsilon}|\Upsilon(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{\Upsilon}^{2}}\frac{\langle 0|J_{\nu}^{J/\psi}|J/\psi(q,\varepsilon_{2})\rangle}{q^{2}-m_{J/\psi}^{2}}$
$\times\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}(p\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
Besides, one should specify the matrix element $\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}(p)\rangle$ which depends on the momenta and polarization vectors of the particles $\mathcal{M}$, $\Upsilon$ and $J/\psi$
@@
$\left.\times\varepsilon_{1}^{\ast}\cdot\varepsilon_{2}^{\ast}-q\cdot\varepsilon_{1}^{\ast}p^{\prime}\cdot\varepsilon_{2}^{\ast}\right].$ (34)
As a result, for $\Pi_{2\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ we get the expression
$\Pi_{2\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=g_{2}(q^{2})\frac{\Lambda f_{\Upsilon}m_{\Upsilon}f_{J/\psi}m_{J/\psi}}{\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{\Upsilon}^{2})(q^{2}-m_{J/\psi}^{2})}$
$\times\left[\frac{(m^{2}-m_{\Upsilon}^{2}-q^{2})}{2}g_{\mu\nu}-p_{\nu}^{\prime}q_{\mu}\right]+\cdots.$ (35)
The function $\Pi_{2\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$, i.e., the QCD component of the sum rule is given by the expression
$\Pi_{2\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ib}(y-x)\right$.
$\left.\times\gamma_{5}S_{c}^{bj}(x)\gamma_{\nu}S_{c}^{ja}(-x)\gamma_{5}S_{b}^{ai}(x-y)\right].$ (36)
@@
$\times e^{m^{2}/M_{1}^{2}}e^{m_{\Upsilon}^{2}/M_{2}^{2}}\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$
@@
for $(M_{2}^{2},s_{0}^{\prime})$ satisfy all these requirements. Note that $s_{0}^{\prime}$ is limited by the mass $m_{\Upsilon(2S)}=(10023.4\pm 0.5)~\mathrm{MeV}\$of the radially excited state $\Upsilon(2S)$, i.e., $s_{0}^{\prime}<m_{\Upsilon(2S)}^{2}$.
@@ III.3 $\mathcal{M}\rightarrow B_{c}^{+}B_{c}^{-}$
$\times J^{B_{c}^{-}}(0)J^{\dagger}(x)\}|0\rangle,$ (42)
@@
$\langle B_{c}^{+}(p^{\prime})B_{c}^{-}(q)|\mathcal{M}\rangle=g_{3}(q^{2})p\cdot p^{\prime},$ (45)
@@
$\times\frac{m^{2}+m_{B_{c}}^{2}-q^{2}}{2(q^{2}-m_{B_{c}}^{2})}+\cdots.$ (46)
@@
$\left.\times\gamma_{5}S_{b}^{ai}(x-y)\right]\mathrm{Tr}\left[\gamma_{5}S_{c}^{bj}(x)\gamma_{5}S_{b}^{jb}(-x)\right].$ (47)
@@
$\times e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}}^{2}/M_{2}^{2}}\Pi_{3}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (48)
@@ IV.1 $\mathcal{M}\rightarrow D^{0}\overline{D}^{0}$ and $D^{+}D^{-}$
$\widetilde{\Pi}_{1}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{D^{0}}(y)$ (52)
$\times J^{\overline{D}^{0}}(0)J^{\dagger}(x)\}|0\rangle.$
@@
First, we find the physical side of the SR and write down $\widetilde{\Pi}_{1}(p,p^{\prime})$ in the following way
$\widetilde{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J^{D^{0}}|D^{0}(p^{\prime})\rangle}{p^{\prime 2}-m_{D^{0}}^{2}}\frac{\langle 0|J^{\overline{D}^{0}}|\overline{D}^{0}(q)\rangle}{q^{2}-m_{D^{0}}^{2}}$
$\times\langle D^{0}(p^{\prime})\overline{D}^{0}(q)|\mathcal{M}(p)\rangle\frac{\langle\mathcal{M}(p)|J^{\dagger}|0\rangle}{p^{2}-m^{2}}+\cdots,$
@@
To simplify $\widetilde{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})$, we rewrite it in terms of the involved particles’ masses and current couplings (decay constants). The matrix element of the pseudoscalar $D^{0}$ ($\overline{D}^{0}$) meson is given by the expression
@@
Afterwards, we recast $\widetilde{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})$ into the following form
$\widetilde{\Pi}_{1}^{\mathrm{Phys}}(p,p^{\prime})=\widetilde{g}_{1}(q^{2})\frac{\Lambda f_{D}^{2}m_{D}^{4}}{2m_{c}^{2}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{D^{0}}^{2}\right)}$
$\times\frac{\left(m^{2}+m_{D^{0}}^{2}-q^{2}\right)}{(q^{2}-m_{D^{0}}^{2})}+\cdots,$ (57)
Note that the whole expression in Eq. (57) is the invariant amplitude $\widetilde{\Pi}_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ which can be employed to get SR for the form factor $\widetilde{g}_{1}(q^{2})$.
@@
$\widetilde{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle\overline{b}b\rangle$
$\times\mathrm{Tr}\left[\gamma_{5}S_{c}^{ja}(y-x){}S_{c}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right],$ (58)
@@
The correlation function $\widetilde{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})$ depends on three propagators and vacuum condensate $\langle\overline{b}b\rangle$ of $b$ quarks. This function differs from standard correlators which contain four heavy quark propagators $S_{b(c)}(x)$. In fact, to calculate $\widetilde{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})$ one contracts heavy and light quark fields, and because mesons $D^{0}\overline{D}^{0}$ do not contain $b$ quarks gets an expression with the condensate $\langle\overline{b}b\rangle$.
For further studies, we make use of the relation between condensates
$\langle\overline{b}b\rangle=-\frac{1}{12m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle$ (59)
derived in Ref. Shifman:1978bx from the sum rule analysis. This expression was obtained at the leading order of the perturbative QCD and is valid provided that the higher order corrections in $m_{b}^{-1}$ are very small. Having activated this equality, we get
$\widetilde{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})=-\frac{1}{36m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
$\times\mathrm{Tr}\left[\gamma_{5}S_{c}^{ja}(y-x){}S_{c}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (60)
Stated differently, the correlator $\widetilde{\Pi}_{1}^{\mathrm{OPE}}(p,p^{\prime})$ is suppressed by the factor $\langle\alpha_{s}G^{2}/\pi\rangle$. We denote the corresponding invariant amplitude as $\widetilde{\Pi}_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ and carry out all required operations to fix SR for the form factor $\widetilde{g}_{1}(q^{2})$.
We omit standard details and present only final results for the process $\mathcal{M}\rightarrow D^{0}\overline{D}^{0}$. Numerical computations of the form factor $\widetilde{g}_{1}(q^{2})$ are performed using